OpenStudy (xguardians):

Square root of 18 minus Square root of 50

OpenStudy (anonymous):

I was showing you the steps, but you keep leaving, which implies you just want the answer.

OpenStudy (xguardians):

Sorry i'm knew to this website? I'm not sure how do you show me the steps? I really need to learn its just a review for me! Sorry about that ~ Thanks though!

OpenStudy (anonymous):

Alright, np, I'll show you it.

OpenStudy (anonymous):

\[\sqrt{18}-\sqrt{50}\] \[\sqrt{18} = \sqrt{2\times3^3} = 3\sqrt{2}\] \[\sqrt{50}=\sqrt{2\times5^2}=5\sqrt{2}\] \[3\sqrt{2}-5\sqrt{2} \implies -2\sqrt{2}\]

OpenStudy (anonymous):

that should be a 3^2 not 3^3

OpenStudy (xguardians):

Thank you very much! Wasn't sure how to approach this problem! Now I know to make common square root radicals :D

OpenStudy (anonymous):

\[\sqrt{2\times3^2}=3\sqrt{2}\]

OpenStudy (anonymous):

Np